### v1

# Online Matroid Embeddings

Abstract
We introduce the notion of an online matroid embedding, which is an algorithm for mapping an unknown matroid that is revealed in an online fashion to a larger-but-known matroid. The existence of such embedding enables a reduction from the version of the matroid secretary problem where the matroid is unknown to the version where the matroid is known in advance. We show that online matroid embeddings exist for binary (and hence graphic) and laminar matroids. We also show a negative result showing that no online matroid embedding exists for the class of all matroids. Finally, we define the notion of an approximate matroid embedding, generalizing the notion of $\alpha$-partition property, and provide an upper bound on the approximability of binary matroids by a partition matroid, matching the lower bound of Dughmi et al.
1 Introduction
A common setup in online algorithms is to have a matroid whose structure is revealed to the algorithm one element at a time. The algorithm processes the elements of the ground set in sequence and at each point in time, it has access to the dependencies between the elements that have already arrived. Typical examples are the matroid secretary problem and matroid prophet inequalities.
In such problems, is there any advantage in knowing the matroid structure in advance? Imagine the following situation: we are processing an unknown matroid M, however, we know a fixed (potentially very large) matroid BigM that has an isomorphic copy of every possible matroid M and moreover, we can construct this embedding online. We will show that if such object exists, then we can reduce the version of the problem where the matroid is revealed online to the version of the problem where the matroid structure is known, by assuming our matroid is BigM.
Moreover, if the on-the-fly embedding maintains uniform random order, then the existence of such an online embedding implies that in that class the matroid secretary problem with unknown structure is no harder than the matroid secretary problem with known structure.
1.1 Two versions of the Matroid Secretary Problem
The Matroid Secretary Problem (MSP) is one of the most exciting open questions in online algorithms and a lot of effort was spent over time to understand many of its variants and special cases. In the MSP a sequence of weighted elements is presented to an algorithm in random order. The goal of the algorithm is to select elements that form an independent set of a certain matroid M and whose combined weight is in expectation an $\alpha$-approximation to the weight of the optimal basis.
In the original definition of the problem by Babaioff et al [5, 6], the algorithm has access to the matroid M in advance. The only information missing is the weight of the elements. The algorithm is presented with the labels of elements in M in random order together with their weight. Certain algorithms like Korula-Pál [22] for graphic matroids and Dinitz-Kortsarz [11] for regular matroids, rely on knowing the structure to pre-compute a decomposition of the matroid which is then used by the online algorithm.
This is in contrast with the $O(\sqrt{\log\textsf{rank}})$-approximation algorithm by Chakraborty-Lachish [9], and the more recent $O(\log\log\textsf{rank})$-approximation algorithms by Lachish [24] and Feldman-Svensson-Zenklusen [15] that only require an independence oracle over the set of elements observed so far.
The main application of the techniques we develop is to study reductions between the two versions of this problem, the known-matroid MSP and the online-revealed-matroid MSP. Before we go on and describe our main contribution, we leave the reader with two open versions of the MSP (one of each type):
MSP for binary matroids and known vectors:
An input of this problem consists of pairs $(u_{i},w_{i})$ where $u_{i}$ is a vector in $\mathbb{F}_{2}^{n}$ and $w_{i}$ is a non-negative weight. The input is presented to the algorithm in random order. The algorithm sees both the vectors and the weights and selects $\mathbb{F}_{2}$-linearly independent vectors.
MSP for graphic matroids and unknown graph:
An input consists of a graph with edge set $E$ and weights on the edges. The input is again presented to the algorithm in random order but the algorithm can only see the labels of the edges ($\textsf{edge}_{1}$, $\textsf{edge}_{2}$, $\textsf{edge}_{3}$, …) and doesn’t know the vertices they connect. Instead, the algorithm has access to an oracle which can answer for a subset of edges if it contains a cycle or not. The algorithm must select a set of edges that are independent.
An interesting consequence of our reduction in particular shows that a solution to the first problem implies a solution with the same approximation ratio for the second problem.
1.2 Our Contribution
1.2.1 Online Matroid Embedding
Our main conceptual contribution is to define the notion of an online matroid embedding (OME) in Section 3. For a given class of matroids $\mathcal{C}$ and a host matroid BigM, we define an OME as a set of matroid monomorphisms, i.e., mappings that preserve the matroid structure, from any matroid in the class $\mathcal{C}$ into BigM that can be constructed sequentially, only using calls to an independence oracle over the set of elements observed so far.
The use of embeddings in algorithm design is an idea that has been successfully explored in other contexts, most notably, metric embeddings both in classic settings [8, 25, 3] as well as more recently in online settings [18, 4, 28]. While our motivation and main application is the matroid secretary problem, we believe that understanding maps between matroids preserving structure is an important mathematical question in its own right that can enable other algorithmic applications beyond the MSP.
1.2.2 Consequence for the MSP
In Theorem 5.1 we show that the existence of an OME for a class of matroids enables a reduction from the online-revealed-matroid MSP to the known-matroid MSP. The challenge in proving this is to show that the online embedding can be used in a way that (almost) maintains uniform random arrival order. We do so by interleaving elements in BigM that are in the image of M with the remaining elements in BigM. This interleaving does not ensure uniformity of the arrival order of the elements in BigM, but it’s close to being so in total variation distance. We complete the reduction with a coupling argument that shows that the existence of an $\alpha$-competitive algorithm for the known-matroid MSP implies the existence of a $(\alpha-\epsilon)$-approximate algorithm for the online-revealed MSP.
Our reduction implies that an algorithm for the MSP over binary matroids cannot meaningfully use any advance information about the matroid. This is in contrast to all known $O(1)$-competitive algorithms for special cases of binary matroids [22, 11].
1.2.3 Constructing OMEs
In Section 4 (Theorem 4.2 and Theorem 4.4) we provide a complete analysis for binary matroids. Namely, for the class of binary matroids M with $n$ elements, there is an online matroid embedding into BigM the complete binary matroid $\mathbb{F}_{2}^{n}$. We also develop a technique for making the OME order-independent: we use properties of the automorphism group of $\mathbb{F}_{2}^{n}$ together with randomization to ensure that the images of the elements in M are not correlated with the arrival order. This technique is in fact more general (Theorem 4.6) and can be applied whenever the group of automorphisms of the host matroid is “sufficiently rich” in a sense that the theorem statement makes precise.
The key property of binary matroids that we exploit to establish these results is that in $\mathbb{F}_{2}^{n}$, there is a unique element that completes a circuit, in the sense that there cannot be two circuits of the same size that intersect in all but one element of each.
We refer to online matroid embeddings where both M and BigM are of the same class as “within-class” OMEs. In Section 6 we explore whether such “within-class” OMEs can exist for graphic matroids. We show that such embeddings cannot exist, in fact we show that graphic matroids cannot be embedded in an online-fashion to regular matroids. To rule out the existence of such an online embedding, we show that if it would exist, then BigM must contain an isomorphic copy of $\mathbb{F}_{2}^{n}$. However, $\mathbb{F}^{2}_{n}$ contains an isomorphic copy of the Fano plane which is not representable over $\mathbb{F}_{3}$ [31]. Hence BigM can’t be regular.
We believe that the lack of online matroid embeddings for graphic matroids/regular matroids that “don’t leave the class” may shed light on why progress on the known-matroid MSP for graphic and regular matroids has not extended to the online-revealed version of these problems, and more generally the MSP for general binary matroids.
In Section 7 we give another example of an OME, namely for laminar matroids. We show (in Theorem 7.3) how to embed the class of laminar matroids M with at most $n$ elements into BigM which is a complete linear matroid of rank $n$ over any field with sufficiently many elements.
Finally in Section 8 we show an impossibility result of constructing an OME for the class of all matroids. This is shown by studying finite projective planes and showing that for those matroids, elements that haven’t arrived yet impose non-trivial constraints on the already arrived elements. As a corollary we obtain an impossibility of constructing an OME for the class of all matroids representable over fields of characteristic at least $7$.
1.2.4 Approximate OMEs
In Section 9 we extend the notion of an OME to allow distortion, i.e., the map approximately preserves the rank. We observe that the $\alpha$-partition property in Abdolazimi et al. [1] and Dughmi et al. [12] can be viewed as an approximate matroid embedding into the free matroid. Combining our formalism with their lower bounds on embedding into the free matroid, we also provide a lower bound on the distortion needed to embed the complete binary matroid into a graphic matroid (Corollary 9.3). This result is an example of the power of the formalism: with the right definitions, extending the lower bound to larger classes becomes a simple corollary. In Theorem 9.6 we show the tightness of the $\Omega(n/\log n)$ lower bound in [12] of the distortion of embedding the complete binary matroid into the free matroid by constructing an embedding achieving this distortion. Finally, we state two open questions on whether we can use distortion to bypass the impossibility results in ealier sections.
1.3 Related Work
Matroid Secretary Problem
The matroid secretary problem was first studied in [5, 6, 7], who gave a $O(\log(\textsf{rank}))$-competitive algorithm for general matroids. This bound was improved to $O(\sqrt{\log(\textsf{rank})})$ in [9], and the state-of-the-art is a $O(\log\log(\textsf{rank}))$-competitive algorithm [24, 15]. The algorithms of [24, 15] only uses independence oracle calls on subsets of the elements revealed so far.
For graphic matroids there is a $O(1)$-competitive algorithm, provided that the graphic matroid is known in advance [22]. The same is true for the more general class of regular matroids [11]. Laminar matroids also admit an $O(1)$-competitive algorithm [20, 21]. Some evidence for the difficulty of the matroid secretary problem for general binary matroids can be found in [26] and [1], showing that binary matroids are not $(b,c)$-decomposable, ruling out a promising approach to obtaining an $O(1)$-competitive algorithm for this class.
Oveis Gharan and Vondrák [29] systematized the study of matroid secretary problem variants, establishing a notation for classifying problem variants according to whether the elements arrive in adversarial or random order, whether the assignment of weights to elements is adversarial or random, and whether or not the matroid structure is known in advance. In their nomenclature, the main question addressed in our work is whether the RO-AA-MK variant is equivalent to the RO-AA-MN variant for matroids in general, or for specific classes of matroids. Interestingly, for variants with adversarial arrival order but random weight assignment, [29] demonstrates a stark qualitative difference in approximability: the AO-RA-MK model (when the matroid structure is known in advance) admits a 64-competitive algorithm for all matroids, whereas the AO-RA-MN model (when the number of elements is known in advance but the matroid structure is revealed online) has no constant-competitive algorithm even for the class of rank one matroids!
Very recently, [30] gave a $O(1)$-competitive algorithm for the matroid secretary problem in the random assignment model when the matroid structure is not known in advance, and instead is only revealed over time.
Matroid Prophet Inequalities
The matroid prophet inequality problem was first studied in [17]. An asymptotically optimal $(1-o(1))$-competitive algorithm for $k$-uniform matroids was given in [2]. A tight $O(1)$-competitive algorithm for the matroid prophet inequality problem was given in [23], also see [10] for the problem of maximizing submodular functions subject to matroid constraints. Constant-factor competitive algorithms can also be obtained via online contention resolution schemes (OCRS) [16]. Random-order versions of the matroid prophet inequality problem are studied in [13].
To the best of our knowledge, all these algorithms exploit that the matroid structure is known in advance. An additional difficulty for reductions of the type we present in this paper, is that typically these algorithms need to know the identity of the distribution that a certain element’s weight is drawn from. For the i.i.d. case this is obviously not an obstacle, and so our reductions apply.
Metric Embeddings and Distortion
An important inspiration for this work comes from the literature on metric embeddings. A classic result in this context is Bourgain’s theorem [8]. The algorithmic importance of such embeddings, and Bourgain’s theorem in particular, was first highlighted in the seminal papers of [25, 3].
Since then metric embeddings have found applications in a host of algorithmic problems, see, e.g., the survey of [19] and Chapter 15 of [27]. Closer to our notion of online matroid embeddings is a recent line of work on online metric embeddings [18, 4, 28] in which points of a metric space are presented one at a time to an algorithm who must then decide on a mapping to the host metric space. The main difference is that instead of preserving a matroid structure, those papers try to minimize metric distortion.
2 Matroids, Morphisms, and $K$-representations
Throughout the paper, we will use $[n]$ to denote the set of integers $\{1,2,\ldots,n\}$.
Matroid Definition
A matroid M is composed by a ground set $M$ and a rank function $\textsf{rank}_{\textbf{M}}:2^{M}\rightarrow\mathbb{Z}_{+}$ satisfying the following properties:
  • $\textsf{rank}_{\textbf{M}}(\emptyset)=0$;
  • $\textsf{rank}_{\textbf{M}}(S\cup\{i\})-\textsf{rank}_{\textbf{M}}(S)\in\{0,1\},\forall S,\{i\}\subseteq M$
  • $\textsf{rank}_{\textbf{M}}(S\cup T)+\textsf{rank}_{\textbf{M}}(S\cap T)\leq\textsf{rank}_{\textbf{M}}(S)+\textsf{rank}_{\textbf{M}}(T),\forall S,T\subseteq M$ (submodularity)
It follows from the second condition that $\textsf{rank}_{\textbf{M}}(S)\leq|S|$. Whenever $|S|=\textsf{rank}_{\textbf{M}}(S)$ we say that $S$ is an independent set of the matroid. Otherwise, we say that $S$ is dependent. A minimal dependent set is called a circuit, i.e., $C\subseteq M$ is a circuit if $C$ is dependent but every strict subset $S\subsetneq C$ is independent. We say that a matroid has rank $r$ if $r=\max_{S}\textsf{rank}_{\textbf{M}}(S)$.
We say that an element $x\in\textbf{M}$ is a loop if $\textsf{rank}_{\textbf{M}}(\{x\})=0$. We say that a matroid is loop-free if every set of one element is independent. Given a set $S\subseteq\textbf{M}$ we define the span as $\textsf{span}_{\textbf{M}}(S)=\{x\in\textbf{M};\textsf{rank}_{\textbf{M}}(S\cup\{x\})=\textsf{rank}_{\textbf{M}}(S)\}$.
Matroid Morphisms
We will use the same notation to refer to a matroid and its ground set. Given two matroids M and N we will define a morphism $f:\textbf{M}\rightarrow\textbf{N}$ to be a map between their ground sets that preserves rank, i.e.:
$\textsf{rank}_{\textbf{N}}(f(S))=\textsf{rank}_{\textbf{M}}(S),\forall S\subseteq\textbf{M}.$
Whenever the matroid morphism is an injective map, we will say it is a matroid monomorphism or a matroid embedding. Whenever it is bijective, we will say it is a matroid isomorphism. An isomorphism from a matroid to itself is called an automorphism. (Aside: this paragraph defines the category of matroids in the sense of category theory. However, we won’t use any other fact from category theory other than borrowing its very convenient language.)
[…]
Element Copies
Given a matroid M and an integer $k$ we will define the matroid $\textbf{M}_{[k]}$ by creating $k$ copies of each element of M. Formally, the ground set of $\textbf{M}_{[k]}$ is $\{(u,j);u\in\textbf{M},j\in[k]\}$. The rank function of $\textbf{M}_{[k]}$ is induced by the projection $\phi:\textbf{M}_{[k]}\rightarrow\textbf{M}$ that maps $(u,j)\mapsto u$, i.e., $\textsf{rank}_{\textbf{M}_{[k]}}(S)=\textsf{rank}_{\textbf{M}}(\phi(S))$. By definition, the projection $\phi$ is a matroid morphism from $_{[k]}\rightarrow\textbf{M}$ M.
If N is a matroid of at most $n$ elements, every morphism $f:\textbf{N}\rightarrow\textbf{M}$ can be written as: $f=\phi\circ f^{\prime}$ where $f^{\prime}:\textbf{N}\rightarrow\textbf{M}_{[n]}$ is a monomorphism.
Direct Sum
[…]
Graphic Matroids
We will define a few special classes of interest. We start with graphic matroids. Given a graph with edge set $E$, we can define a matroid with ground set $E$ by defining the $\textsf{rank}(S)$ of a subset $S\subseteq E$ as the maximum number of edges in $S$ that don’t form a cycle. We say that a matroid M is graphic if it is isomorphic to the matroid obtained from an undirected graph as we just described.
As an example, consider the matroid M with ground set $\{a,b,c\}$ and rank function such that $\textsf{rank}(S)=|S|$. The matroid is graphic since it is isomorphic to the matroid that can be obtained from any of the graphs in Figure 1. An important thing to note, however, is that the matroid description contains no information about vertices. It only tells us which sets of edges are independent and which are not. As we can see in the figure, this is typically not enough to fully determine the graph structure.
$a$$b$$c$$a$$b$$c$
Figure 1: Two graphs that generate the same matroid on their edge set
$K$-representable Matroids
Let $K$ be a field (e.g. $\mathbb{Q},\mathbb{R},\mathbb{F}_{p}$) and let $K^{d}$ be the vector space formed by $d$-dimensional vectors with coordinates in $K$. We say that a subset of vectors $u_{1},\ldots,u_{k}\in K^{d}$ is independent if the unique solution to $\alpha_{1}u_{1}+\alpha_{2}u_{2}+\ldots+\alpha_{k}u_{k}=0$ for $\alpha_{i}\in K$ is $\alpha_{1}=\alpha_{2}=\ldots=\alpha_{k}=0$. Any subset of $K^{d}$ together with the independency relation above defines a matroid. From now on, we will use the notation $K^{d}$ to represent both the vector space and the corresponding matroid. We say that a matroid M is $K$-representable if there is a matroid morphism $\textbf{M}\rightarrow K^{d}$ for some integer $d$.
If a matroid M is $K$-representable for every field $K$ we say that M is a regular matroid. Every graphic matroid is representable over any field by mapping an edge $(u,v)$ to the vector $e_{u}-e_{v}$ where $e_{u}$ the the $u$-th unit vector. (This is true even over $\mathbb{F}_{2}$ where $e_{u}-e_{v}=e_{u}+e_{v}$.)
For example, the matroids in Figure 1 can be represented by the vectors $(1,-1,0,0)$, $(0,1,-1,0)$, $(0,0,1,-1)$. As it is the case for graphic matroids, the matroid description has no information about vectors and the representation is again not unique. An equally good representation is $(1,0,0),(0,1,0),(0,0,1)$.
We will be specially interested in binary matroids which are matroids that are representable over $\mathbb{F}_{2}$ (the finite field of $2$ elements where addition and multiplication are performed mod $2$).
Laminar matroids
A family of sets, $\mathcal{A}$, is called laminar if it satisfies the property that for any $A,A^{\prime}\in\mathcal{A}$, at least one of the sets $A\cap A^{\prime},\,A\setminus A^{\prime},\,A^{\prime}\setminus A$ is empty. A laminar matroid M is one for which there exists a laminar family of sets $\mathcal{A}$ consisting of subsets of the ground set of M and a function $c:\mathcal{A}\to\mathbb{Z}_{+}$, such that the independent sets of M are precisely those sets $I\subseteq M$ such that $|I\cap A|\leq c(A)$ for all $A\in\mathcal{A}$.
Uniform Matroid
We will denote by $\textbf{U}_{n,r}$ the uniform matroid of $n$ elements and rank $r$. This is the matroid with ground set $[n]$ and whose rank function is $\textsf{rank}_{\textbf{U}_{n,r}}(S)=\min(r,|S|)$. We call the $\textbf{Fr}_{n}:=\textbf{U}_{n,n}$ the free matroid of rank $n$, i.e., a matroid of $n$ elements in which every element is independent.
Trivial Matroid
Let T be the trivial matroid which has ground set $\{0\}$ and rank function $\textsf{rank}_{\textbf{T}}(S)=0$ for all sets $S$.
3 Online Matroid Embeddings
We are interested in studying matroids whose structure is revealed to an algorithm in an online fashion. For that, it will be useful to take into account the order in which elements are processed, which we will represent by an indexing of the ground set: $\pi:[n]\rightarrow\textbf{M}$.
A matroid N is a restriction of matroid M if the ground set of N is a subset of the ground set of M and $\textsf{rank}_{\textbf{M}}$ coincides with $\textsf{rank}_{\textbf{N}}$ on the ground set of N.
Given a matroid with ordered ground set specified by a pair $(\textbf{M},\pi)$, we say that $(\textbf{M}^{\prime},\pi^{\prime})$ is a prefix-restriction of $(\textbf{M},\pi)$ if $n^{\prime}=|\textbf{M}^{\prime}|<|\textbf{M}|$, $\textbf{M}^{\prime}$ is the restriction of M to $\pi([n^{\prime}])$ and $\pi^{\prime}$ is the restriction of $\pi$ to $[n^{\prime}]$.
Let $\mathcal{C}$ be a class of matroids that is closed under restriction (e.g., the class of all matroids, graphic matroids, binary matroids, $K$-representable matroids, matroids of rank at most $r$). An online matroid morphism (OMM) for class $\mathcal{C}$ consists of a host matroid BigM, together with matroid morphisms
$f_{\textbf{M},\pi}:\textbf{M}\rightarrow\textbf{BigM}$
for every $\textbf{M}\in\mathcal{C}$ and every indexing $\pi:[n]\rightarrow\textbf{M}$ of the ground set of M, such for every prefix-restriction $(\textbf{M}^{\prime},\pi^{\prime})$ of $(\textbf{M},\pi)$, the map $f_{\textbf{M}^{\prime},\pi^{\prime}}$ is the restriction of $f_{\textbf{M},\pi}$ to the ground set of $\textbf{M}^{\prime}$.
If all morphisms $f_{\textbf{M},\pi}$ are monomorphisms, we say that they form an online matroid embedding (OME). Given an online matroid morphism it is easy to construct an online matroid embedding by copying the elements of BigM.
Lemma 3.1.
Let $\mathcal{C}$ be a class of matroids, where each matroid $\textbf{M}\in\mathcal{C}$ has at most $n$ elements and $f_{\textbf{M},\pi}:\textbf{M}\rightarrow\textbf{BigM}$ form an online matroid morphism. Then there is an online matroid embedding $f^{\prime}_{\textbf{M},\pi}:\textbf{M}\rightarrow\textbf{BigM}_{[n]}$.
Proof.
We define $f^{\prime}_{\textbf{M},\pi}$ as follows: for each $u\in\textbf{M}$ if $u=\pi(k)$ let $f^{\prime}_{\textbf{M},\pi}(u)=(f_{\textbf{M},\pi}(u),j)$ where $j=|i\in[k];f_{\textbf{M},\pi}(\pi(i))=u|$. The functions $f^{\prime}_{\textbf{M},\pi}$ are injective by construction and they are matroid morphisms by the definition of $\textbf{BigM}_{[n]}$. In fact: $f_{\textbf{M},\pi}=\phi\circ f^{\prime}_{\textbf{M},\pi}$ where $\phi$ is the natural projection $_{[n]}\rightarrow\textbf{BigM}$ BigM. Finally note that they can be constructed online since the identity of the copy used is only a function of the set of elements that arrived up to this point. ∎
With this definition we can ensure that an online algorithm is able to construct a monomorphism from an unknown matroid M to BigM in an online fashion. Consider a matroid M for which the elements arrive according to $\pi$. At each time $t$, we can observe the structure of the matroid $\textbf{M}_{t}$ which is the restriction of M to $\pi([t])$. Let $\pi_{t}$ be the restriction of $\pi$ to $[t]$. If we have an online matroid embedding, we can first construct $f_{\textbf{M}_{1},\pi_{1}}$, then extend to $f_{\textbf{M}_{2},\pi_{2}}$ and so forth.
It will also be convenient to define a randomized online matroid morphism (embedding) which for every matroid $\textbf{M}\in\mathcal{C}$ and ordering $\pi$ specifies a distribution over (mono)morphisms $f_{\textbf{M},\pi}:\textbf{M}\rightarrow\textbf{BigM}$ such that for every prefix-restriction $(\textbf{M}^{\prime},\pi^{\prime})$ the distribution of the restriction of $f_{\textbf{M},\pi}$ to the ground set of $\textbf{M}^{\prime}$ coincides with the distribution of $f_{\textbf{M}^{\prime},\pi^{\prime}}$.
Finally, we say that a randomized online matroid embedding is order-independent if the distribution of $f_{\textbf{M},\pi}$ doesn’t depend on $\pi$. In other words, for any two orderings $\pi$ and $\pi^{\prime}$, the morphisms $f_{\textbf{M},\pi}$ and $f_{\textbf{M},\pi^{\prime}}$ are equally distributed.
Online vs. Offline Embeddings
The difficulty of constructing an online matroid embedding is that the elements of BigM corresponding to certain elements of M must be chosen before the full matroid structure of M is known. If we merely wanted to construct a matroid BigM that contains an isomorphic copy of every matroid in $\mathcal{C}$, that would be very easy: BigM could be taken to be the direct sum of all the matroids in $\mathcal{C}$.
4 Binary Matroids
Before we discuss how to use online matroid embeddings in online algorithms, it is important to show first that they exist in non-trivial cases. For that, we will provide a complete analysis for binary matroids. Recall that $\mathbb{F}_{2}^{n}$ is the complete binary matroid of rank $n$ and that graphic matroids and regular matroids are special cases of binary matroids. Our first result is the existence of an OME for this class. This will be done by showing the existence of an OMM and using Lemma 3.1 to convert an OMM to an OME.
Our first step is to show a lemma that the matroid $\mathbb{F}_{2}^{n}$ is special in the sense that its group of matroid automorphisms coincides with its group of vector space automorphisms:
Lemma 4.1.
A mapping $A:\mathbb{F}_{2}^{n}\rightarrow\mathbb{F}_{2}^{n}$ is a matroid automorphism iff it is an automorphism of vector spaces.
Proof.
An automorphism of vector spaces $A:\mathbb{F}_{2}^{n}\rightarrow\mathbb{F}_{2}^{n}$ is a bijection such that for any vectors $v_{1},\ldots,v_{k}\in\mathbb{F}_{2}^{n}$ it holds that $A(\sum_{i=1}^{k}v_{i})=\sum_{i=1}^{k}Av_{i}$. This in particular implies that a set of vectors $v_{1},\ldots,v_{k}$ is independent iff the vectors $Av_{1},\ldots,Av_{k}$ are independent. This is because a coefficient vector $(\alpha_{1},\ldots,\alpha_{k})$ satisfies the equation $\sum_{i=1}^{k}\alpha_{i}v_{i}=0$ if and only if it satisfies $\sum_{i=1}^{k}\alpha_{i}Av_{i}=A(\sum_{i=1}^{k}\alpha_{i}v_{i})=0$, so the first equation has only trivial solutions if and only if the second equation has only trivial solutions.
For the opposite direction, if $A$ is a matroid automorphism and $e_{1},\ldots,e_{n}$ is the standard basis of $\mathbb{F}_{2}^{n}$ then $Ae_{1},\ldots,Ae_{n}$ must be linearly independent elements of $\mathbb{F}_{2}^{n}$. Now, take any vector $v=\sum_{i\in S}e_{i}$. Since $\{v\}\cup\{e_{i};i\in S\}$ forms a circuit, then $\{Av\}\cup\{Ae_{i};i\in S\}$ must form a circuit. Since the only non-zero constant in $\mathbb{F}_{2}$ is $1$, it must hold that: $Av+\sum_{i\in S}Ae_{i}=0$ and hence $Av=\sum_{i\in S}Ae_{i}$. Hence $A$ is also an automorphism of vector spaces. ∎
Theorem 4.2.
Let $\mathcal{C}$ be the class of binary matroids of at most $n$ elements and let BigM be the complete binary matroid $\mathbb{F}_{2}^{n}$. Then there exists an OMM for $\mathcal{C}$ into BigM.
Proof.
Given a binary matroid M we construct a mapping $f:\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$ as follows. We keep a counter $k$ initially set to $1$. For each element $a$ we process, if it is independent of the previously arrived elements (i.e. there are no circuits containing $a$ and the elements seen so far), we set $f(a)=e_{k}$ and increment $k$. Otherwise, $a$ forms a circuit with a set of previously arrived elements $u_{1},\ldots,u_{m}$ for some integer $m\geq 0$. This means that their image $f(a),f(u_{1}),\ldots,f(u_{m})$ must be a minimal $\mathbb{F}_{2}$-linearly dependent set. Since the only non-zero constant in $\mathbb{F}_{2}^{n}$ is $1$, then it must hold that:
$f(a)+f(u_{1})+\ldots+f(u_{m})=0$
and hence we can map: $f(a)$ to $f(u_{1})+\ldots+f(u_{m})$ (recall that $1=-1$ in $\mathbb{F}_{2}$).
Finally, we need to argue that $f$ is a matroid morphism. Observe that if M is a binary matroid, then there exists a morphism $g:\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$. Let $\{b_{1},\ldots,b_{r}\}$ be the elements of M such that $f(b_{i})=e_{i}$. By the fact that $g$ is matroid morphism, $g(b_{1}),\ldots,g(b_{r})$ are linearly independent elements in $\mathbb{F}_{2}^{n}$. By Lemma 4.1 there is an automorphism $A\in\text{Aut}(\mathbb{F}_{2}^{n})$ that takes $g(b_{i})$ to $e_{i}$. Since matroid morphisms compose, $Ag:\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$ is matroid morphism.
[…]
Finally, we argue that $f(a)=Ag(a)$ for all $a$ in M. We show this by induction. For each element processed by the algorithm, if it is independent from previously arrived elements, then $f(a)=Ag(a)$ by construction. Otherwise, there are previously arrived elements such that $a,u_{1},\ldots,u_{m}$ form a circuit. Hence their image under $Ag$ must be a linearly dependent set of $\mathbb{F}_{2}$-vectors, which means that:
$Ag(a)=Ag(u_{1})+\ldots+Ag(u_{m})=f(u_{1})+\ldots+f(u_{m})=f(a)$
where the second equality holds by induction. Since $f$ coincides with $Ag$, $f$ is a matroid morphism. ∎
Furthermore, there is a randomized OME that is order-independent. We will show it as a consequence of the following lemmas:
Lemma 4.3.
Given a binary matroid M and the complete binary matroid $\mathbb{F}_{2}^{n}$, if there are two matroid morphisms $f,g:\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$, then there exist an automorphism $A\in\text{Aut}(\mathbb{F}_{2}^{n})$ such that $f=A\circ g$.
Proof.
Let $r$ be the rank of M and let $\{b_{1},\ldots,b_{r}\}$ be a basis of M. Then $\{f(b_{1}),f(b_{2}),\ldots,f(b_{r})\}$ and $\{g(b_{1}),g(b_{2}),\ldots,g(b_{r})\}$ are both sets of independent vectors in $\mathbb{F}_{2}^{n}$. Then there exists an automorphism $A$ of vector spaces (and hence a matroid automorphism) that sends $g(b_{i})$ to $f(b_{i})$. For any other element in $v\in\textbf{M}$ consider any circuit formed with a subset of the basis. If $\{v\}\cup\{b_{i};i\in S\}$ is a circuit then it must be the case that: $f(v)=\sum_{i\in S}f(b_{i})$ and $g(v)=\sum_{i\in S}g(b_{i})$. Given that $A$ is a automorphism of vector spaces, then: $Ag(v)=\sum_{i\in S}Ag(b_{i})=\sum_{i\in S}f(b_{i})=f(v)$. ∎
Theorem 4.4.
There is a order-independent randomized OMM from the class of binary matroids $\mathcal{C}$ into the complete binary matroid.
Proof.
Let $f_{\textbf{M},\pi}$ be the online matroid morphism constructed in Theorem 4.2 and consider $A\circ f_{\textbf{M},\pi}$ when $A$ is drawn uniformly at random from $\text{Aut}(\mathbb{F}_{2}^{n})$. It is clear that for every fixed $A$ the morphisms $A\circ f_{\textbf{M},\pi}$ still form an OMM. We only need to check that they are order-independent. To see that, observe that if $\pi$ and $\pi^{\prime}$ are two different orderings of the ground set of M then $f_{\textbf{M},\pi}$ and $f_{\textbf{M},\pi^{\prime}}$ are two morphisms $\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$. By the previous lemma, there is $A_{0}\in\text{Aut}(\mathbb{F}_{2}^{n})$ such that $f_{\textbf{M},\pi}=A_{0}\circ f_{\textbf{M},\pi^{\prime}}$. Now, the distribution of $A\circ f_{\textbf{M},\pi}$ for a random $A\sim\text{Aut}(\mathbb{F}_{2}^{n})$ is the same distribution as $A\circ A_{0}\circ f_{\textbf{M},\pi^{\prime}}$ which is the same distribution of $A\circ f_{\textbf{M},\pi^{\prime}}$, since $A_{0}\circ A$ is also uniformly distributed over $\text{Aut}(\mathbb{F}_{2}^{n})$. ∎
Extending to Copies
In the following section we will be needing an online matroid embedding. For that reason, we need to extend the last two theorems to deal with copies. The extension is rather simple: we only need to observe that a matroid autormorphism of the commplete binary matroid with $n$ copies of each element $(\mathbb{F}_{2}^{n})_{[n]}$ can be decomposed into an automorphism $A\in\text{Aut}(\mathbb{F}_{2}^{n})$ and indexings of the identities of the copies.
Lemma 4.5.
If $f\in\text{Aut}((\mathbb{F}_{2}^{n})_{[n]})$ then there exists $A\in\text{Aut}(\mathbb{F}_{2}^{n})$ and indexings $\sigma_{u}:[n]\rightarrow[n]$ for each $u\in\mathbb{F}_{2}^{n}$ such that $f((u,j))=(Au,\sigma_{u}(j))$
Proof.
Let $id$ be the identity map and $\phi:(\mathbb{F}_{2}^{n})_{[n]}\rightarrow\mathbb{F}_{2}^{n}$ the natural projection. Now, $\phi\circ f$ and $\phi$ are two matroid morphisms from $(\mathbb{F}_{2}^{n})_{[n]}\rightarrow\mathbb{F}_{2}^{n}$ so by Theorem 4.3 there is $A\in\text{Aut}(\mathbb{F}_{2}^{n})$ such that $\phi\circ f=A\circ\phi$ (see the commutative diagram below). This means in particular that $f((u,j))=(Au,\sigma_{u}(j))$ for some indexings $\sigma_{u}$.
${(\mathbb{F}_{2}^{n})_{[n]}}$${\mathbb{F}_{2}^{n}}$${(\mathbb{F}_{2}^{n})_{[n]}}$${(\mathbb{F}_{2}^{n})_{[n]}}$${\mathbb{F}_{2}^{n}}$$\scriptstyle{\phi}$$\scriptstyle{id}$$\scriptstyle{f}$$\scriptstyle{\phi}$$\scriptstyle{A}$
With that, Theorem 4.4 automatically extends to the matroid with copies $(\mathbb{F}_{2}^{n})_{[n]}$ by taking a random automorphism from group $\text{Aut}((\mathbb{F}_{2}^{n})_{[n]})$.
Remark A (Single Orbit Morphisms)
[…]
Theorem 4.6 (Generalization of Theorem 4.4).
Let $f_{\textbf{M},\pi}$ be an OMM of class $\mathcal{C}$ into BigM such that given two orderings $\pi$ and $\pi^{\prime}$ of the ground set of M, there is an automorphism $A\in\text{Aut}(\textbf{BigM})$ such that $f_{\textbf{M},\pi^{\prime}}=A\circ f_{\textbf{M},\pi}$. Then there is an order independent randomized OMM of class $\mathcal{C}$ into BigM.
Remark B (Other Fields)
In this section we repeatedly use the fact that $\mathbb{F}_{2}$ has only one non-zero constant, so whenever we identify a circuit in the matroid, we know exactly what is the linear dependency between the elements in the corresponding vector field. This is no longer true even in slightly larger fields like $\mathbb{F}_{3}$. If vectors $u,v,w\in\mathbb{F}_{3}^{n}$ form a circuit, it could be that: $w=\pm u\pm v$ in the representation. As a consequence, given two matroid morphisms $f,g:\textbf{M}\rightarrow\mathbb{F}_{3}^{n}$ there may not exist a vector-space automorphism $A$ of $\mathbb{F}_{3}^{n}$ such that $f=Ag$. In the previous example, if $f(u)=g(u)$, $f(v)=g(v)$ but $f(w)=f(u)+f(v)$ but $g(w)=g(u)-g(v)$ no such automorphism can exist.
5 OMEs and the Matroid Secretary Problem
We consider two versions of the matroid secretary problem (MSP):
  • Known-matroid MSP: In the first version, a matroid M is known to the algorithm ahead of time. The only information missing is the weight of each element. The algorithm then processes pairs $(u,w_{u})$ in random order, where $u$ is an element in the ground set of M, and $w_{u}\in\mathbb{R}_{+}$ is its weight. Upon seeing the element, the algorithm must irrevocably decide whether to accept that element or not, subject to the constraint that the set of accepted elements must be an independent set of M.
  • Online-revealed-matroid MSP: In the second version, the algorithm has only access to the number of elements $n=|\textbf{M}|$ and to a promise that $\textbf{M}\in\mathcal{C}$ for a class of matroids $\mathcal{C}$. The algorithm again processes pairs $(u,w_{u})$ in random order at each time, but it only knows the rank function restricted to the subset of elements that have already arrived.
The first of these is identified as the RO-AA-MK model in the nomenclature of [29], whereas the second is identified as the RO-AA-MN model. The second version is clearly harder in the sense that an $\alpha$-approximation to the second version immediately implies an $\alpha$-approximation to the first version. Our goal is to understand when is the converse true. We will show the following theorem:
Theorem 5.1.
If a class $\mathcal{C}$ of matroids admits a randomized order-independent online matroid embedding into matroid BigM, then an $\alpha$-approximation to the Known-matroid MSP for $\textbf{BigM}_{[n]}$ implies that for every $\epsilon>0$ there is a $(\alpha-\epsilon)$-approximation to the Online-revealed-matroid MSP for $\mathcal{C}$.
For the case of binary matroids that we previously discussed, the matroids BigM and $\textbf{BigM}_{[n]}$ themselves are binary in which case we can obtain the following corollary:
Corollary 5.2.
For binary matroids, there is no gap in approximability of the known-matroid MSP and the online-revealed-matroid MSP.
The rest of this section is dedicated to a reduction from the known-matroid MSP for $\textbf{BigM}_{[n]}$ together with an online matroid embedding to the online-revealed-matroid MSP. The main challenge of the reduction is to (mostly) preserve random order of arrivals.
5.1 Proof of Theorem 5.1
Let M be the unknown matroid that is revealed to the algorithm in an online fashion and let $f_{\textbf{M},\pi}$ be an order-independent randomized OME into BigM. Let $n=|\textbf{M}|$, $N=|\textbf{BigM}|$ and $d,k$ be two larger integers (to be specified later) where $k$ is a multiple of $d$.
An instance of the MSP consists of a sequence of weighted elements from M that are presented to the algorithm in random order. Our goal is to map it on the fly to a random instance of the MSP on $\textbf{BigM}_{[k]}$. The main difficulty is, as usual, doing it online and preserving the random order. Our strategy will be to first provide an offline reduction which preserves random order and obtains the desired approximation, but can’t be implemented online. After that we will provide a mostly-online implementation of this reduction, i.e. a procedure that samples from the same distribution generated by the offline reduction and that with $1-\epsilon$ probability can be implemented online. With the remaining $\epsilon$ probability, the process raises a flag. Raising a flag will indicate that from that point on, the reduction can no longer be implemented online. Algorithmically, we will stop the algorithm whenever we raise a flag and obtain zero reward. Finally, we will show that the probability of raising a flag is very small for large values of $k$ and $d$.
Offline Reduction
We will view a weighted element of M as a pair $(u,w_{u})$ with $u\in\textbf{M}$ and $w_{u}\in\mathbb{R}_{+}$. In the offline reduction, we assume we have access to the entire matroid M and the entire sequence of weights. Now, we will produce a distribution of instances of $\textbf{BigM}_{[k]}$ as follows.
For each element $v$ in BigM sample $k$ different i.i.d. timestamps $t_{vj}$ for $j\in[k]$ from the $\text{Uniform}([0,1])$ distribution. Those timestamps specify the arrival time of each of the $k$ copies of the elements in BigM and induce a random ordering over the ground set of $\textbf{BigM}_{[k]}$. For the matroid M, sample a random embedding $f:\textbf{M}\rightarrow\textbf{BigM}$ from the OME. (Since the embedding is order independent, we don’t need to know the arrival order of elements in M to sample such embedding). For each $u\in\textbf{M}$, pick a random copy of $f(u)$ and set its weight to $w_{u}$. For the remaining elements, set the weight equal to zero.
This random input is clearly in random order as it is equivalent to starting with $k$ copies of the elements of BigM where all but one copy has weight zero if that corresponds to an element of M and randomly permuting those elements. Now, feed this instance to the $\alpha$-competitive algorithm for the MSP on $\textbf{BigM}_{[k]}$. From the set selected by the algorithm, discard any element with zero weight chosen by the algorithm. The elements with non-zero weight chosen in BigM correspond to an independent set of M with the same weight. Hence, in expectation, we select an $\alpha$-approximation to the optimal basis of M.
Mostly-online Implementation
The drawback of the previous reduction is that it can’t be implemented online as we are assuming we know everything in advance. We will describe the same sampling procedure in a way that with high probability we can generate the instance as we go. In the sampling procedure, we will also define an event raise a flag which will mean that we can’t generate that instance online as we learn the structure of the matroid M.
The process will again start by sampling i.i.d. timestamps $t_{vj}$ for $v\in\textbf{BigM}$ and $j\in[k]$ from $\text{Uniform}([0,1])$. In addition, we will also sample $n$ additional timestamps from $\text{Uniform}([0,1])$ sort them in increasing order and denote the sorted list by $T_{1},\ldots,T_{n}$.
We will now divide the interval $[0,1]$ into intervals $I_{i}=[\frac{i-1}{d},\frac{i}{d})$ for $i\in[d]$. If more than one timestamp $T_{s}$ falls in the same interval $I_{i}$, we will raise a flag. We will count how many of the timestamps $t_{v,j}$ fall in each interval:
$X_{vi}=|\{j\in[k];t_{vj}\in I_{i}\}|$
With that, also define:
$A_{vi}=\min\left(\frac{k}{d},X_{vi}\right)\qquad B_{vi}=\max\left(0,X_{vi}-\frac{k}{d}\right)$
Now, process the elements of the matroid M according to order $\pi$ (which will be sampled at random). As we process the $s$-th element $u=\pi(s)\in\textbf{M}$, we will map it to an element in $v=f_{\textbf{M},\pi}(u)\in\textbf{BigM}$ using the randomized OME. Now, we will apply the following procedure to choose a copy of $v$ in $\textbf{BigM}_{[k]}$ to assign weight $w_{u}$:
  • find the interval $I_{i}$ containing $T_{s}$.
  • with probability $A_{vi}d/k$, choose one of the $X_{vi}$ timestamps $t_{vj}$ in interval $I_{i}$
  • with remaining probability (if any), raise a flag and choose a different interval $I_{i^{\prime}}$ with probability proportional to $B_{vi^{\prime}}$ and choose a timestamp $t_{vj}$ in that interval.
We assign weight $w_{u}$ to the element with the chosen timestamp and zero weight to others. Now, we will show the following facts.
Lemma 5.3.
The mostly online implementation samples sequences with the same probability as the offline reduction.
Proof.
Observe that if we ignore the weights, the order of the elements of BigM is the same in both processes since they are determined by the timestamps $t_{vj}$. What we are left to argue is that for each $v$ we select uniformly random timestamp $t_{vj}$ to assign the non-negative weight. For that, observe that since $f_{\textbf{M},\pi}$ is order independent, it has the same distribution as if we first sampled a monomorphism $f:\textbf{M}\rightarrow\textbf{BigM}$, and then we sampled an independent uniform indexing $\pi$ for the arrival order of the elements in $f(\textbf{M})$. This implies that when we assign the timestamps $T_{1},\dots,T_{n}$ according to $\pi$, the resulting distribution is the same as if we assigned i.i.d. Uniform$[0,1]$ timestamps $T_{v}$ to each element $v\in f(\textbf{M})$, and therefore, the interval $T_{v}$ lands in is uniformly chosen and independent across elements $v\in f(\textbf{M})$. Now fix a certain timestamp $t_{vj}$ and let $I_{i}$ be the interval containing it. We will show that the probability that this timestamp is selected is exactly $1/k$.
Consider two cases: either $X_{vi}\leq k/d$ in which case the probability of sampling $t_{vj}$ is the probability that the timestamp $T_{v}$ is in $I_{i}$ times (which is $1/d$), times the probability we decide to sample a timestamp inside $I_{i}$ (which is $X_{vi}d/k$), times the probability that out of those, we choose $t_{vj}$ (which is $1/X_{vi}$). The total probability is:
$\frac{1}{d}\cdot\frac{X_{vi}d}{k}\cdot\frac{1}{X_{vi}}=\frac{1}{k}$
In the case where $X_{vi}>k/d$, then it is possible that we sample $t_{vj}$ also when $T_{v}$ is outside $I_{i}$. The probability that we sample $t_{vj}$ and $T_{v}$ is in $I_{i}$ is:
$\frac{1}{d}\cdot 1\cdot\frac{1}{X_{vi}}=\frac{1}{dX_{iv}}$
The probability that we sample when it is outside is the probability that we choose a different interval $I_{i^{\prime}}$, raise a flag and then move to interval $I_{i}$, which is:
$\sum_{i^{\prime}}\frac{1}{d}\left(1-\frac{A_{vi^{\prime}}d}{k}\right)\cdot\frac{B_{vi}}{\sum_{i^{\prime}}B_{vi^{\prime}}}\cdot\frac{1}{X_{vi}}=\frac{1}{k}\frac{\left(k-\sum_{i^{\prime}}A_{vi^{\prime}}\right)}{\sum_{i^{\prime}}B_{vi^{\prime}}}\cdot\frac{B_{vi}}{X_{vi}}=\frac{1}{k}\frac{B_{vi}}{X_{vi}}$
because $\sum_{i^{\prime}}A_{vi^{\prime}}+\sum_{i^{\prime}}B_{vi^{\prime}}=\sum_{i^{\prime}}X_{vi^{\prime}}=k$. Taking those two probabilities together, we have:
$\frac{1}{dX_{iv}}+\frac{1}{k}\frac{B_{vi}}{X_{vi}}=\frac{1}{k}.\qed$
Lemma 5.4.
If no flags were raised, we can produce the instance on the fly as we process M.
Proof.
Let $i_{1}<\ldots<i_{n}$ be the indices of the intervals such that $T_{s}\in I_{i_{s}}$. Since no flag was raised, then each $T_{s}$ landed in a different interval and the $s$-th element that arrives from matroid M is mapped to a copy inside $I_{i_{s}}$. This enables the following online reduction: once the $s$-th element arrives we can decide the weights of all the elements in intervals $I_{i_{s-1}+1}$ to $I_{i_{s}}$ and feed to the MSP algorithm for $\textbf{BigM}_{[k]}$. In this sub-sequence there will be at most one element on non-zero weight which corresponds to the arriving element of M. We can observe if that element was selected in $\textbf{BigM}_{[k]}$ and if so, we can select it in M. ∎
Lemma 5.5.
For any $n$ and $\epsilon$, there are large enough $k$ and $d$, such that the probability that we raise a flag is at most $\epsilon$.
Proof.
The first event in which we raise a flag is when two timestamps $T_{s}$ land in the same interval. The probability that this happens is at most $n^{2}/d$. Now, note that for each interval $i$ and each of the $n$ elements $v$ in BigM that have non-zero weights, we have by the Chernoff bound that:
$\mathbb{P}\left(X_{vi}\leq(1-\delta)\frac{k}{d}\right)\leq\exp\left(-\frac{\delta^{2}k}{2d}\right)$
Hence with probability at most $nd\exp\left(-\frac{\delta^{2}k}{2d}\right)$, the timestamps $t_{vj}$ are such that the probability we raise a flag when we try to choose a timestamp in the same interval as $T_{s}$ is more than $\delta n$. Taking the union bound of those events, we get:
$\frac{n^{2}}{d}+nd\exp\left(-\frac{\delta^{2}k}{2d}\right)+\delta n$
Taking $\delta=\epsilon/(3n)$, $d=3n^{2}/\epsilon$ and $k$ large enough, we get that the total probability of raising a flag is at most $\epsilon$. ∎
Taking those lemmas together, we can conclude the proof of Theorem 5.1. For that, let Alg be an $\alpha$-competitive algorithm for $\textbf{BigM}_{[k]}$ and let $Y$ represent the sequence of the MSP sampled by the offline reduction. Let’s represent by $\textsf{Alg}(Y)$ the weight of the elements selected by Alg and Opt the weight of the optimal basis. By the fact that the offline reduction produces an instance in random order, we know that $\mathbb{E}[\textsf{Alg}(Y)]\geq\alpha\textsf{Opt}$.
Our online reduction, will attempt to construct $Y$ on the fly. If we raise the flag, we will stop the algorithm and pretend we had zero reward. If not, we will continue the reduction and collect $\textsf{Alg}(Y)$ reward. We will denote by Flag the event that the flag was raised and by $\overline{\textsf{Flag}}$ its complement. Our total reward will be:
$\mathbb{E}[\textsf{Alg}(Y)\cdot\mathbf{1}\{\overline{\textsf{Flag}}\}]=\mathbb{E}[\textsf{Alg}(Y)]-\mathbb{E}[\textsf{Alg}(Y)\cdot\mathbf{1}\{\textsf{Flag}\}]\geq\mathbb{E}[\textsf{Alg}(Y)]-\textsf{Opt}\cdot\mathbb{P}[\textsf{Flag}]\geq(\alpha-\epsilon)\textsf{Opt}.$
6 Graphic and Regular Matroids
For graphic matroids, there exists an elegant $2e$-approximation algorithm for the MSP in the known-matroid case by Korula-Pál [22]. Their algorithm assumes that when an edge arrives, the algorithm learns the pair of vertices it connects as well as the weight. In other words, the algorithm processes in each step $(u,v),w_{uv}$ where $u$ and $v$ are vertices. Even though the full graph is not known in advance, there is enough information about the graph structure to randomly decompose the problem into instances of the single-item secretary problem.
In the online-revealed matroid case, the algorithm has only access to an oracle that tells which subsets of previously arrived edges are independent (contain no cycles). For example, if $3$ edges arrive and the algorithm knows that they are all independent, it is impossible for the algorithm to know if they form a path, a star or if they share not endpoints (recall Figure 1). Korula-Pál heavily relies on having vertex information and doesn’t easily extend to this model. In fact, we are not aware of any $O(1)$-approximation algorithm for the MSP for graphic matroids in the online-revealed-matroid setting.
In the remainder of this section we investigate whether we can obtain such an algorithm using a OME. A natural idea is to try to construct an embedding where the host matroid BigM is itself a graphic matroid. If one could do that, it would be possible to combine the reduction in Section 5 with the Korula-Pál algorithm to obtain an algorihtm for the graphic MSP in the online-revealed-matroid setting. Unfortunately, we show below that no such embedding exists:
Theorem 6.1.
If $\mathcal{C}$ is the class of graphic matroids, there is no online matroid embedding into a host matroid BigM where BigM is also graphic.
Proof.
Assume that such embedding exists and let $G=(V,E)$ be the graph representing BigM. Now, consider two graphic matroids $\textbf{M}_{1}$ and $\textbf{M}_{2}$ represented respectively by the graphs in the left and right of Figure 2. Observe that when restricted to $\{a,b,c\}$, the matroids are identical since every non-empty subset of elements is independent. When an OMM observes the restriction to those three elements, it has not way to know in which matroid we are in, so it needs to map those three edges to the same edges of the graph $G$ representing BigM. Let’s denote those edges by $a^{\prime},b^{\prime},c^{\prime}$. Let also $d^{\prime},e^{\prime},f^{\prime}$ be the edges in $G$ that the edges of the left matroid are mapped to and let $g^{\prime}$ be the edge that the $g$ edge in the right matroid is mapped to. Note that $\{a^{\prime},b^{\prime},d^{\prime}\}$, $\{b^{\prime},c^{\prime},e^{\prime}\}$ and $\{a^{\prime},c^{\prime},f^{\prime}\}$ must form cycles in $G$. The only way that this is possible while keeping $\{a^{\prime},b^{\prime},c^{\prime}\}$ independent is if edges $a^{\prime},b^{\prime},c^{\prime}$ all share an endpoint. However, edges $a^{\prime},b^{\prime},c^{\prime},g^{\prime}$ must form a cycle in $G$ as well, which is not possible if the first three edges share an endpoint. ∎
$a$$b$$c$$d$$f$$e$$a$$b$$c$$g$
Figure 2: Two graphic matroids whose restriction to $\{a,b,c\}$ coincide.
Graphic matroids are a special case of regular matroids, for which there exists an $O(1)$-competitive algorihtm by Dinitz-Kortsarz [11]. It is then tempting to construct an embedding from graphic into a host matroid BigM that is regular. However, that is again not possible:
Lemma 6.2.
If a host matroid BigM admits an online matroid embedding of all rank $n$ graphic matroids, then BigM must also admit an online matroid embedding of all rank $n$ binary matroids.
Proof.
We will show that BigM contains an isomorphic copy of $\mathbb{F}_{2}^{n}$. We first define for each $S\subseteq[n]$ a graphic $\textbf{M}_{S}$ with ground set $[n+1]$ represented by the graph where the only cycle is formed by the edges with labels in $S\cup\{n+1\}$. Equivalently, the rank function is given by
$\textsf{rank}_{\textbf{M}_{S}}(T)=\left\{\begin{aligned} &|T|-1&&\text{for }T\supseteq S\cup\{n+1\}\\ &|T|&&\text{otherwise }\end{aligned}\right.$
Let also $\textbf{M}^{\prime}$ be the matroid on $[n]$ such that $\textsf{rank}_{\textbf{M}^{\prime}}(T)=|T|$ for all subsets $T$. Assuming all ground serts are ordered according to the labels, note that $\textbf{M}^{\prime}$ is a prefix of all matroids $\textbf{M}_{S}$. Now, let $b_{1},\ldots,b_{n}\in\textbf{BigM}$ be the image of the ground set of $\textbf{M}^{\prime}$ by the online embedding. Since BigM is regular, it admits an $\mathbb{F}_{2}$-representation, so we can think of $b_{1},\ldots,b_{n}$ as linearly independent vectors in $\mathbb{F}_{2}^{n}$.
Since this is an online embedding, it can be extended to a morphism $\textbf{M}_{S}\rightarrow\textbf{BigM}$ for each $S\subseteq[n]$. So, for each $S$, let $b_{S}$ be the element in BigM $[n+1]$ maps to. If we view $b_{S}$ as an $\mathbb{F}_{2}$-vector, we must have $b_{S}=\sum_{i\in S}b_{i}$ in $\mathbb{F}_{2}^{n}$. As a consequence, the vectors $\{b_{S};S\subseteq[n]\}$ form an isomorphic copy of $\mathbb{F}_{2}^{n}$. ∎
We can now derive the following theorem as as corollary:
Theorem 6.3.
If $\mathcal{C}$ is the class of graphic matroids, there is no online matroid embedding into a host matroid BigM where BigM is regular.
Proof.
By the previous lemma, if such embedding exists then BigM must contain an isomorphic copy of $\mathbb{F}_{2}^{n}$. However, $\mathbb{F}_{2}^{n}$ contains an isomorphic copy of the Fano plane which is not representable over $\mathbb{F}_{3}$ [31]. Hence BigM can’t be regular. ∎
The reader will notice that Theorem 6.1 can be derived as a trivial corollary of Theorem 6.3 since graphic matroids are regular. Nevertheless, we find the more direct proof of Theorem 6.1 enlightening and opted to keep it.
7 Laminar Matroids
Another important class of matroids that admits an online embedding is the class of laminar matroids. We begin this section by recalling the definition of a laminar matroid and proving a useful structural lemma about them. We then present an OMM for the class of laminar matroids with at most $n$ elements, using a host matroid BigM which is a complete linear matroid of rank $n$ over any field with sufficiently many elements.
Lemma 7.1.
If $f:\textbf{M}\to\textbf{N}$ is a matroid monomorphism and N is laminar, then M is laminar as well.
Proof.
[…]
$\tilde{c}(B)=\min\{c(A);A\in\mathcal{A}\mbox{ and }B=f^{-1}(A)\}.\qed$
Let us also recall the definitions of the span and flat in a matroid. Given a subset $S$ of the ground set of a matroid M, we define the $\textsf{span}_{\textbf{M}}(S)=\{x\in\textbf{M};\textsf{rank}_{\textbf{M}}(S\cup\{x\})=\textsf{rank}_{\textbf{M}}(S)\}$. From the properties of the rank function, this is the equivalent to the maximal set containing $S$ that has the same rank as $S$. We say that a subset $S$ is a flat if $S=\textsf{span}(S)$. For the complete $K$-representable matroid $K^{n}$, the span coincides with the notion of the linear span of vector fields and the flats are linear subspaces.
The span of laminar matroids has the following useful property:
Theorem 7.2 (Fife and Oxley [14]).
A matroid is laminar if and only if, for all circuits $C_{1}$ and $C_{2}$ with $C_{1}\cap C_{2}\neq\emptyset$, either $\textsf{span}(C_{1})\subseteq\textsf{span}(C_{2})$ or $\textsf{span}(C_{2})\subseteq\textsf{span}(C_{1})$.
Theorem 7.3.
Let $\mathcal{C}$ be the class of laminar matroids of at most $n$ elements and $\mathbb{F}$ be a field with at least $2^{n}$ elements. Then there is an OMM from $\mathcal{C}$ into $\mathbb{F}^{n}$.
Proof.
The OMM is defined inductively. If M is an empty matroid then $f_{\textbf{M},\pi}$ is the trivial morphism from the empty matroid to $\textbf{BigM}=\mathbb{F}^{n}$. Otherwise let $n=|\textbf{M}|$, let $(\textbf{M}^{\prime},\pi^{\prime})$ denote the restriction of $(\textbf{M},\pi)$ to $[n-1]$, let $g$ denote the morphism $f_{\textbf{M}^{\prime},\pi^{\prime}}$, and let $u=\pi(n)$. The morphism $f=f_{\textbf{M},\pi}$ is defined as follows. For $u^{\prime}\neq u$ we set $f(u^{\prime})=g(u^{\prime})$. If $u$ doesn’t belong to any circuit with previously arrived elements, define $f(u)$ to be any element outside the linear span of $f(\textbf{M}^{\prime})$. If $u$ is in some circuit, define $A\subseteq\textbf{M}$ to be the intersection of $\textsf{span}(C)$ over all circuits $C$ in M containing $u$. By Theorem 7.2, $A$ is the span of some circuit in M and hence a flat. Now, choose $f(u)$ to be an element in the linear subspace $V=\textsf{span}_{\mathbb{F}^{n}}(f(A\setminus\{u\}))$ that is not contained in $V_{S}=\textsf{span}_{\mathbb{F}^{n}}(f(S))$ for every subset $S$ of M such that $S\cup\{u\}$ is independent on M.
To justify that $f(u)$ is well-defined, we must argue that the linear subspace $V$ contains at least one element that is not in $V_{S}$ for every subset $S$ such that $S\cup\{u\}$ is independent. We will use a counting argument that consists of showing that $V_{S}\cap V$ is a proper linear subspace of $V$ for every such $S$, and then observing that a vector space over a field with at least $2^{n}$ elements cannot be expressed as the union of $2^{n}$ or fewer proper linear subspaces. To show that $V_{S}\cap V$ is a proper linear subspace of $V$ we argue by contradiction. Let $C$ be a circuit containing $u$ such that $A=\textsf{span}_{\textbf{M}}(C)$. As $S\cup\{u\}$ is independent, there must be some $u^{\prime}\in C\setminus\{u\}$ such that $S\cup\{u^{\prime}\}$ is also independent. Then, since $g:\textbf{M}^{\prime}\to\textbf{BigM}$ is a matroid morphism, $g(u^{\prime})$ is linearly independent of $g(S)$. In particular, $g(u^{\prime})\in V\setminus V_{S}$ cannot and hence $V_{S}\cap V$ is a proper linear subspace of $V$ as claimed.
Having justified that $f$ is well-defined, we must show that it is a matroid morphism. Consider any set $S$ in M. To show $\textsf{rank}_{\textbf{BigM}}(f(S))=\textsf{rank}_{\textbf{M}}(S)$ we will proceed by case analysis.
  1. 1.
    If $u\not\in S$ then $f(S)=g(S)$ and we use the induction hypothesis that $g=f_{\textbf{M}^{\prime},\pi^{\prime}}$ is a matroid morphism.
  2. 2.
    If $u\in S$ and $\textsf{rank}_{\textbf{M}}(S)=\textsf{rank}_{\textbf{M}}(S\setminus\{u\})+1$ then let $B$ be a maximal independent subset of $S\setminus\{u\}$. Since $g$ is a matroid morphism, every $u^{\prime}\in S\setminus\{u\}$ lies in the linear span of $g(B)$. Since $B\cup\{u\}$ is independent in M, by construction $f(u)$ lies outside the linear span of $g(B)$. Hence, $\textsf{rank}_{\textbf{BigM}}(S)=\textsf{rank}_{\textbf{BigM}}(B\cup\{u\})=|B|+1=\textsf{rank}_{\textbf{M}}(S)$.
  3. 3.
    If $u\in S$ and $\textsf{rank}_{\textbf{M}}(S)=\textsf{rank}_{\textbf{M}}(S\setminus\{u\})$ then there is a circuit $C\subseteq S$ containing $u$. The $\textsf{span}(C\setminus\{u\})$ is a flat $A^{\prime}$ that contains $u$, so $A^{\prime}$ must be a superset of $A$ by Theorem 7.2. Then we have the following chain of containments.
    $\textsf{span}_{\mathbb{F}^{n}}(g((S\setminus\{u\}))\supseteq\textsf{span}_{\mathbb{F}^{n}}(g((C\setminus\{u\}))=\textsf{span}_{\mathbb{F}^{n}}(g((A^{\prime}))\supseteq\textsf{span}_{\mathbb{F}^{n}}(g(A\setminus\{u\})).$
    Since $f(u)$, by construction, belongs to $\textsf{span}_{\mathbb{F}^{n}}(g(A\setminus\{u\}))$, it belongs to $\textsf{span}_{\mathbb{F}^{n}}(g(S\setminus\{u\}))$ and therefore
    $\textsf{rank}_{\mathbb{F}^{n}}(f(S))=\textsf{rank}_{\mathbb{F}^{n}}(g(S\setminus\{u\}))=\textsf{rank}_{\textbf{M}^{\prime}}(S\setminus\{u\})=\textsf{rank}_{\textbf{M}}(S).\qed$
Similarly to the situation of graphic matroids, we can embed laminar matroids into a large linear matroid, but there is no embedding for which BigM is also laminar:
Theorem 7.4.
If $\mathcal{C}$ is the class of laminar matroids of at most $n$ elements, there is no online matroid embedding into a host matroid BigM where BigM itself is laminar.
Proof.
Given a set $S$ let $\textbf{U}_{n,r}(S)$ denote the uniform matroid with $n$ elements and rank $r$ defined on ground set $S$. Using this notation, consider the following three matroids on ground set $\{a,b,c,d\}$:
$\textbf{U}_{3,2}(\{a,b,d\})\oplus\textbf{U}_{1,1}(\{c\})\qquad\textbf{U}_{3,2}(\{a,c,d\})\oplus\textbf{U}_{1,1}(\{b\})\qquad\textbf{U}_{3,2}(\{b,c,d\})\oplus\textbf{U}_{1,1}(\{a\})$
It is easy to check that all three of those matroids are laminar. Moreover, the restriction to $\{a,b,c\}$ is the matroid $\textbf{U}_{3,3}(\{a,b,c\})$. Assume now that there is an OMM into a laminar matroid BigM and let $a^{\prime},b^{\prime},c^{\prime}$ be the elements it maps to. There must be some set in the laminar family of BigM that separates those elements, otherwise we can’t extend it to the three matroids above. For example, in the first matroid we can’t swap the roles of $a$ and $c$ preserving rank. Finally, there is at most one two-element subset of $\{a^{\prime},b^{\prime},c^{\prime}\}$ that can be formed by intersecting one of the sets in the laminar family with $\{a^{\prime},b^{\prime},c^{\prime}\}$. If the two-element subset that spans $d$ is not that one, then you have no way of extending the embedding to include $d$. ∎
8 Matroids of Bounded Rank
Finally, we show that there is no universal host matroid BigM such that all matroids of at most $n$ elements admit an OMM into BigM. In fact, we show that even if we restrict to matroids of rank $3$ this is not possible. The geometric intuition is that once we reach rank $3$, we start being able to represent finite projective planes, where elements that haven’t arrived yet impose non-trivial constraints on the already arrived elements. Before we get there, it is useful to analyze ranks $1$ and $2$.
Let $\mathcal{M}_{n,r}$ be the matroids of rank at most $r$ defined on at most $n$ elements. We will show that for $r=1,2$ it is trivial to construct an online matroid embedding. As usual, we will construct an OMM and convert it to an OME using Lemma 3.1.
Lemma 8.1.
There is an online matroid morphism from $\mathcal{M}_{n,1}$ into $\textbf{U}_{1,1}\oplus T$.
Proof.
The matroid $\textbf{U}_{1,1}\oplus T$ has ground set $\{0,1\}$ and the only independent set is $\{1\}$. Give a matroid $\textbf{M}\in\mathcal{M}_{n,1}$ and an element $e$ of the ground set of M we map it to $1$ if $\textsf{rank}_{\textbf{M}}(\{e\})=1$ and to 00 otherwise. The mapping is constructed online since it only depends on the rank of a single-set element consisting of the element we are processing. It is also simple to check that it is a morphism since M has rank $1$. ∎
Lemma 8.2.
There is an online matroid morphism from $\mathcal{M}_{n,2}$ into $\textbf{U}_{n,2}\oplus T$.
Proof.
The matroid $\textbf{U}_{n,2}\oplus T$ has ground set $\{0,1,2,\ldots,n\}$ and the independent sets are subsets of at most $2$ elements that don’t contain zero. To construct an embedding $f:\textbf{M}\rightarrow\textbf{U}_{n,2}\oplus T$, when we process an element $e$, we first check if $\textsf{rank}_{\textbf{M}}(\{e\})=0$. If so, we map it to 00. Otherwise, for every element $a$ processed before $e$, we check if $\textsf{rank}_{\textbf{M}}(\{a,e\})=1$. If so, we map $f(e)=f(a)$. Otherwise, we map $e$ to the first unused index in $[n]$. ∎
Once we reach rank $3$, the situation becomes a lot more interesting, as there exist many non-trivial matroids like finite projective planes (e.g. Fano plane). The richness of the space of rank $3$ matroids will also imply that an online matroid embedding no longer exists.
Before we prove it formally, let’s give some geometric intuition. Given points on the plane $\mathbb{R}^{2}$, we can define a matroid of rank $3$ as follows: (i) every set of one point is independent (ii) every pair of different points is independent; (iii) every triple of points is independent iff it they are not collinear; (iv) no other set is independent.
Now, consider the following online problem: we are presented with labels $a,b,c,d,e,f,g$ in this order. Each label represents a point, but we are not told which point it is. Instead, we are told the dependency relation between them. As they arrive, we are asked to map each label to a $\mathbb{R}^{2}$ such that the dependency relations are satisfied.
Take the following instance of this problem: the first $6$ points $a,b,c,d,e,f$ arrive, we are told that all triples are independent, i.e., neither of them is collinear. Figure 3 shows two possible ways to place those points in the plane satisfying those dependencies. In the first arrangement, the segments $[a,b],[c,d],[e,f]$ all intersect in a single point. In the second arrangement, they don’t.
Now, the last label $g$ arrives. If the algorithm chose the arrangement on the left, we can ask to place $g$ such that $\{a,g,b\},\{c,g,d\}$ are dependent, but $\{e,f,g\}$ are independent. There is no way to satisfy those dependencies in the figure on the left, since $g$ must be in the intersection of the $[a,b]$ and $[c,d]$ but the only way to do so is by also being in the line $[e,f]$. On the other hand, if the algorithm chose the arrangement on the right in Figure 3, we can ask to place $g$ such that $\{a,g,b\},\{c,g,d\}$ and $\{e,f,g\}$ are dependent. There is no way to satisfy those dependencies in the figure on the right since the lines $[a,b]$, $[c,d]$ and $[e,f]$ must meet on the same point.
Those two dependencies are satisfiable offline, as we can see in Figure 4. The difficulty is that the dependencies on points that have not arrived pose constraints on the relative position of points that have already arrived.
$a$$e$$c$$b$$f$$d$$a$$e$$c$$b$$f$$d$
Figure 3: Two sets of points that induce the same matroid on $\{a,b,c,d,e,f\}$
$g$$a$$e$$c$$b$$f$$d$$g$$a$$e$$c$$b$$f$$d$
Figure 4: Once the dependency of point $g$ with respect the remaining points is specified, this constrains the possible geometric arrangements of points $\{a,b,c,d,e,f\}$.
Our goal with the previous discussion is to provide an intuition for the following proof. Note that while embedding a matroid of rank $3$, we are not restricted to $\mathbb{R}^{2}$. We could in principle embed it in a matroid that is not representable over any field. However, geometric intuition can now be turned into a combinatorial proof of the following statement:
Theorem 8.3.
There is no host matroid BigM for which there is an online matroid morphism from $\mathcal{M}_{n,3}$ into BigM.
Proof.
Define matroids $\textbf{M}_{1}$ and $\textbf{M}_{2}$ on elements $\{a,b,c,d,e,f,g\}$ represented by the points in Figure 4. In those pictures, a triple of elements is independent iff the corresponding points are not collinear.
Those points induce the same matroid on the first $6$ points, which is simply the uniform matroid $\textbf{U}_{6,3}$, but the matroids differ when we one considers element $g$.
  • in $\textbf{M}_{1}$, the sets $\{a,b,g\}$, $\{c,d,g\}$ and $\{e,f,g\}$ are dependent
  • in $\textbf{M}_{2}$, the sets $\{a,b,g\}$ and $\{c,d,g\}$ are dependent but $\{e,f,g\}$ is independent
Now suppose BigM is a matroid and we’re trying to construct an online embedding of a matroid M into BigM, where M is a rank-3 matroid that could either be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. When the first six elements of M arrive, we have no way of distinguishing whether the input sequence is going to be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. The online embedding algorithm chooses some function $h$ mapping $\{a,b,c,d,e,f\}$ to a six-element subset of BigM. Denote the images of $a,b,c,d,e,f$ in BigM by capital letters, for example $h(a)=A$. Now suppose there are two different extensions of $h$ to the domain ${a,b,c,d,e,f,g}$, denoted by $h_{1}$ and $h_{2}$, such that $h_{i}$ is an embedding of $\textbf{M}_{i}$ into BigM for each $i$. Let $G_{1}=h_{1}(g)$ and $G_{2}=h_{2}(G)$. Then the following must hold:
  • the sets $\{A,B,G_{1}\}$, $\{C,D,G_{1}\}$ and $\{E,F,G_{1}\}$ are dependent in BigM
  • the sets $\{A,B,G_{2}\}$ and $\{C,D,G_{2}\}$ are dependent but $\{E,F,G_{2}\}$ is independent in BigM
The sets $\{A,B,G1,G2\}$ and $\{C,D,G1,G2\}$ both have rank $2$ in BigM, whereas their union has rank $3$ since it contains the rank-3 set $\{A,B,C,D\}$. By submodularity, the set $\{G1,G2\}$ must have rank 1. Again by submodularity,
$\textsf{rank}_{\textbf{BigM}}(\{E,F,G1,G2\})\leq\textsf{rank}_{\textbf{BigM}}(\{E,F,G1\})+\textsf{rank}_{\textbf{BigM}}(\{G1,G2\})-\textsf{rank}_{\textbf{BigM}}(\{G1\})$
which evalutes to $2+1-1=2$. This contradicts the fact that $\{E,F,G2\}$ has rank $3$. ∎
Corollary 8.4.
For every field $K$ of characteristic $p\geq 7$, there is no host matroid BigM for which there is an online matroid embedding from all $K$-representable matroids into BigM.
Proof.
Observe that the matroids $\textbf{M}_{1}$ and $\textbf{M}_{2}$ in the proof of Theorem 8.3 are representable in every field of characteristic at least $7$. For example, they can be represented respectively by the columns of the following matrices:
$\begin{bmatrix}0&2&1&1&2&0&1\\ 1&1&2&0&0&2&1\\ 2&0&0&2&1&1&1\end{bmatrix}\quad\text{ and }\quad\begin{bmatrix}0&2&1&1&2&0&1\\ 1&1&2&0&0&1&1\\ 2&0&0&2&1&1&1\end{bmatrix}$
We know that because of the results in Section 4, Corollary 8.4 is false for $\mathbb{F}_{2}$. We leave it as an open question whether there is an online matroid embedding for class of $\mathbb{F}_{3}$ and $\mathbb{F}_{5}$-representable matroids.
9 Approximate Embeddings
Like metric embeddings, it is useful to extend the notion of matroid embedding to allow distortion. We define an (offline) $\beta$-approximate embedding $f:\textbf{M}\rightarrow\textbf{N}$ as a map between ground sets that approximately preserves rank:
$\frac{1}{\beta}\cdot\textsf{rank}_{\textbf{M}}(S)\leq\textsf{rank}_{\textbf{N}}(f(S))\leq\textsf{rank}_{\textbf{M}}(S),\forall S\subseteq\textbf{M}$
We also define a randomized $\beta$-approximate matroid embedding as a family of functions $f_{r}:\textbf{M}\rightarrow\textbf{N}_{r}$ indexed by a random variable $r\sim R$ such:
$\textsf{rank}_{\textbf{N}_{r}}(f_{r}(S))\leq\textsf{rank}_{\textbf{M}}(S)\text{ a.s. }\forall S\subseteq\textbf{M}$ (1)
$\mathbb{E}[\textsf{rank}_{\textbf{N}_{r}}(f_{r}(S))]\geq\frac{1}{\beta}\textsf{rank}_{\textbf{M}}(S),\forall S\subseteq\textbf{M}$ (2)
Their online counterparts can be defined in the natural way: given a class of matroids $\mathcal{C}$ and host matroid BigM then an online $\beta$-approximate randomized embedding is a family of functions: $f_{\textbf{M},\pi,r}:\textbf{M}\rightarrow\textbf{BigM}$ defined for each $\textbf{M}\in\mathcal{C}$, $\pi:\textbf{M}\rightarrow[n]$ and $r\in R$ such that: (i) it is a randomized $\beta$-approximate matroid embedding for each fixed $\textbf{M},\pi$; (ii) satisfied the prefix-restriction property defined in Section 3.
For example, if we can design a randomized order-independent $\beta$-approximate embedding for class $\mathcal{C}$ into a matroid BigM and there is a known $\alpha$-competitive algorithm for the known-matroid MSP on BigM, then we can use the reduction in Section 5 to convert it into an $(\alpha\beta)$-competitive algorithm for the online-revealed-matroid MSP on class $\mathcal{C}$.
It is worth noting that every loop-free matroid M of rank $n$ admits a trivial $n$-approximate embedding into the free matroid $\textbf{Fr}_{1}:=U_{1,1}$. In the next paragraph, we observe below that we can’t obtain better than $n$ for any graphic host matroid. This is related to the notion called partition property.
Offline Embedding and the Partition Property
The notion of the $\alpha$-partition property was defined by Abdolazimi et al [1] to generalize a property exploited by Korula-Pál [22] in their algorithm for the graphic matroid secretary problem. Translating it to our notation, we say that a matroid M satisfies the $\alpha$-partition property if there is an (offline) $\alpha$-approximate randomized embedding into a free matroid (i.e., a matroid where every non-empty subset is independent). The name partition property comes from the fact that partition matroids are the class of matroids that admit an (exact) morphism to a free matroid. Their respective lower and upper bounds translate to the following results:
Theorem 9.1 (Korula-Pal [22]).
If M is a graphic matroid, then it admits a $2$-approximate randomized embedding into a free matroid.
Theorem 9.2 (Abdolazimi et al[1], Dughmi et al [12]).
If M is the complete binary matroid of rank $n$ and if it admits an $\beta$-approximate randomized embedding into a free matroid, then $\beta\geq\Omega(n/\log n)$. Moreover, there exist a linear matroid M of rank $n$ such that $\beta$-approximate randomized embedding into a free matroid only exist for $\beta\geq\Omega(n)$.
Since composing an $\beta$-approximate embedding with a $\alpha$-approximate embedding we obtain an $(\alpha\beta)$-approximate embedding, we can strenghten the lower bound to also allow for embedding into graphic matroids.
Corollary 9.3.
If M is the complete binary matroid of rank $n$ and if it admits an $\beta$-approximate randomized embedding into a graphic matroid, then $\beta\geq\Omega(n/\log n)$. Moreover, if there is a linear matroid M such that it admits a $\beta$-approximate randomized embedding into a graphic matroid, then $\beta=\Omega(n)$.
As we discussed, constructing an $n$-approximate embedding is trivial. For binary matroids, we complement the lower bound of Dughmi et al [12] with an algorithm to construct a $O(n/\log n)$ embedding into the free matroid. Moreover, this embedding can be computed online.
Matching upper bound that is also online computable
We now construct a randomized approximate embedding from a binary matroid to $\textbf{Fr}_{n}$. Using Theorem 4.2 we can identify every received element $x\in\textbf{M}$ with a vector in $\mathbb{F}_{2}^{n}\setminus\{0\}$. To map $x$ to a partition matroid, we sample a random basis $b_{1},\ldots,b_{n}$ of $\mathbb{F}_{2}^{n}$ and define the function $f_{b}:\mathbb{F}_{2}^{n}\setminus\{0\}\rightarrow[n]$ that maps each $x\in\mathbb{F}_{2}^{n}$ to the smallest index $i$ such that $b_{i}$ is in the unique circuit in $\{x,b_{1},\ldots,b_{n}\}$.
We first show that this embedding doesn’t increase the rank (first property of an approximate embedding in equation (1)):
Lemma 9.4.
Given any loop-free matroid M of rank $n$ and a basis $b_{1},\ldots,b_{n}$ of M, let $f(x)$ be the smallest index $i$ such that $b_{i}$ is in the unique circuit formed by $\{x,b_{1},\ldots,b_{n}\}$. The function $f_{b}:\textbf{M}\rightarrow\textbf{Fr}_{n}$ is such that $\textsf{rank}_{\textbf{Fr}}(f_{b}(S))\leq\textsf{rank}_{\textbf{M}}(S),\forall S\subseteq\textbf{M}$.
Proof.
Given $S$, choose $S^{\prime}\subseteq S$ be such that $\textsf{rank}_{\textbf{Fr}}(f_{b}(S))=\textsf{rank}_{\textbf{Fr}}(f_{b}(S^{\prime}))=|S^{\prime}|$. We will show that $S^{\prime}$ is independent in M and hence $|S^{\prime}|=\textsf{rank}_{\textbf{M}}(S^{\prime})\leq\textsf{rank}_{\textbf{M}}(S)$.
To show that, observe that for any element $x\in\textbf{M}$, $f_{b}(x)=j$ iff $x\in\textsf{span}_{\textbf{M}}(\{b_{j},b_{j+1},\ldots,b_{n}\})\setminus\textsf{span}_{\textbf{M}}(\{b_{j+1},\ldots,b_{n}\})$. If that is the case, then: $\textsf{span}_{\textbf{M}}(\{x,b_{j+1},\ldots,b_{n}\})=\textsf{span}_{\textbf{M}}(\{b_{j},b_{j+1},\ldots,b_{n}\})$.
Now, construct $x_{1},\ldots,x_{n}$ such that $x_{j}$ is the element in $S^{\prime}$ that maps to $j$ if there is such element and $x_{j}=b_{j}$ otherwise. Then $f_{b}(x_{j})=j$. We can show by induction that $\textsf{span}_{\textbf{M}}(\{x_{j},\ldots,x_{n}\})=\textsf{span}_{\textbf{M}}(\{b_{j},\ldots,b_{n}\})$, by replacing elements of $b_{1},\ldots,b_{n}$ one by one by their corresponding $x_{j}$ element and observing that the spans are preserved. In particular, $\{x_{1},\ldots,x_{n}\}$ must be independent since they span the entire matroid. As a consequence the original set $S^{\prime}$ is independent. ∎
We now show a lower bound on the expected rank of $f_{b}(S)$ which will lead to the second property of an approximate embedding:
Lemma 9.5.
For the complete binary matroid $\mathbb{F}_{2}^{n}$, the embedding $f_{b}$ for a random basis $b_{1},\ldots,b_{n}$ of $\mathbb{F}_{2}^{n}$ satisfies:
$\mathbb{E}[\textsf{rank}_{\textbf{Fr}}(f_{b}(S))]\geq(1-1/e)\log_{2}(\textsf{rank}_{\textbf{M}}(S))-O(1),\forall S\subseteq[n]$
Proof.
Step 1: Sampling the basis We consider the following procedure for sampling an uniformly random basis $b_{1},\ldots,b_{n}$ of $\mathbb{F}_{2}^{n}$. For each $i$ we sample an uniform independent vector from $\mathbb{F}_{2}^{n}$ and call it $b^{\prime}_{i}$. Now, if $b^{\prime}_{i}$ is not in $\textsf{span}(b_{1},\ldots,b_{i-1})$ then we set $b^{\prime}_{i}=b_{i}$. Otherwise we resample until we get an element not in the span of the previous elements and set it as $b_{i}$. Nevertheless, we still record $b^{\prime}_{i}$ as the first element sampled such that $b^{\prime}_{1},\ldots,b^{\prime}_{n}$ are i.i.d. vectors. Now, let’s define $\mathcal{E}$ as the event that $b_{i}=b^{\prime}_{i}$ for $i\leq n/2$. We observe that:
$\mathbb{P}(\mathcal{E})=(1-2^{-n})(1-2^{-n+1})\ldots(1-2^{-n/2-1})\geq 1-\textstyle\sum_{k=1}^{n/2}2^{-n+k-1}\geq 1-2^{-n/2}$
Finally, define $B$ to be the $n\times n$ matrix over $\mathbb{F}_{2}$ that has $b_{i}$ as the $i$-th column. Since $b_{1},\ldots,b_{n}$ is a basis, $B$ is invertible.
Step 2: Change of basis Fix a set $S\subseteq\mathbb{F}_{2}^{n}$ such that $r=\min(\textsf{rank}_{\mathbb{F}_{2}^{n}}(S),n/2)$ and let $S^{\prime}=\{s_{1},\ldots,s_{r}\}$ be an independent subset of $S$. Our goal is to bound $\mathbb{E}[\textsf{rank}_{\textbf{Fr}}(f_{b}(S^{\prime}))].$
Let also $X$ be a $n\times n$ invertible matrix over $\mathbb{F}_{2}$ where the first $r$ columns correspond to $s_{1},\ldots,s_{r}$. Define now $\tilde{b}_{i}=XB^{-1}e_{i}$ (where $e_{i}$ is the standard basis). Since the columns of $B$ are uniformly random and $X$ is invertible, $\tilde{b}_{1},\ldots,\tilde{b}_{n}$ is also a uniformly random basis. With that we observe that:
$\mathbb{E}[\textsf{rank}(f_{\tilde{b}_{1}..\tilde{b}_{n}}(S))]\geq\mathbb{E}[\textsf{rank}(f_{\tilde{b}_{1}..\tilde{b}_{n}}(S^{\prime}))]=\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b_{1},\ldots,b_{r}\}))]$
since $b_{i}=BX^{-1}s_{i}$, $e_{i}=BX^{-1}\tilde{b}_{i}$ and $z\mapsto BX^{-1}z$ is a matroid isomorphism of $\mathbb{F}_{2}^{n}$.
Step 3: Conditioning on $\mathcal{E}$ The vectors $b_{1},\ldots,b_{r}$ are not sampled i.i.d. but the vectors $b^{\prime}_{1},\ldots,b^{\prime}_{r}$ are and conditioned on $\mathcal{E}$ they are the same. So we write:
$\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b_{1},\ldots,b_{r}\}))]$ $\geq\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b_{1},\ldots,b_{r}\}))\mid\mathcal{E}]\cdot\mathbb{P}[\mathcal{E}]$
$=\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{r}\}))\mid\mathcal{E}]\cdot\mathbb{P}[\mathcal{E}]$
$\geq\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{r}\}))]-n/2^{n/2}$
Step 4: Bounding the rank in the free matroid The rank in the free matroid is simply the number of elements in the image, so we can re-write:
$\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{r}\}))]$ $=\sum_{j=1}^{n}\mathbb{P}[j\in f_{e_{1}..e_{n}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{r}\}))]$
$=\sum_{j=1}^{n}\left(1-\prod_{i=1}^{r}\mathbb{P}[j\neq f_{e_{1}..e_{n}}(b^{\prime}_{i})]\right)$
$=\sum_{j=1}^{n}\left(1-(1-2^{-j})^{r}\right)$
since $\mathbb{P}[j\neq f_{e_{1}..e_{n}}(b^{\prime}_{i})]$ is the probability that the first non-zero entry of the random vector $b^{\prime}_{i}$ is not $j$, which happens with probability $(1-2^{-j})$. Finally observe that for $j\leq\log_{2}(r)$ we have $(1-2^{-j})^{r}\leq(1-1/r)^{r}\leq 1/e$. Putting this together we obtain that:
$\mathbb{E}[\textsf{rank}(f_{e_{1}..e_{n}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{r}\}))]\geq(1-1/e)\log_{2}(r)\geq(1-1/e)\log_{2}(\textsf{rank}_{\mathbb{F}_{2}^{n}}(S)/2)$
Theorem 9.6.
For any loop-free binary matroid M there is a $O(n/\log n)$-online approximate embedding to the partition matroid.
Proof.
For a loop-free binary matroid M use the composition of the embedding $f_{b}$ for a random basis $b$ of $\mathbb{F}_{2}^{n}$ from the previous lemma with the online morphism $\textbf{M}\rightarrow\mathbb{F}_{2}^{n}$ in Theorem 4.2. Since both can be computed online, their composition can also be computed online.
By Lemma 9.4 the embedding satisfies the first property of an approximate embedding. By Lemma 9.5 a set of $S$ of rank $r$ in M is mapped to a set with expected rank $\max(1,c\log(r)-c^{\prime})$ for constants $c,c^{\prime}$. Finally, observe that $r/\max(1,c\log(r)-c^{\prime})\leq O(n/\log(n))$. ∎
Open Questions.
We end with two open questions on approximate embeddings. The first is whether we can generalize the approximate embedding of Korula-Pál to the online setting. A possible approach is through the construction in Lemma 9.4. The embedding of Korula-Pál is a randomization between two embeddings. The first is exactly the embedding in Lemma 9.4 when the basis $b$ of the graphic matroid is a simple path in the underlying graph. The second can be obtained from a construction analogous to Lemma 9.4 taking the largest index instead of the smallest index. The challenge converting it to the online setting is identifying a ”good basis” in an online fashion knowing only the matroid structure but not the graph.
Open Question 1.
Is there any $O(1)$-approximate online embedding from graphic matroids to the free matroid?
We know that the trivial $n$-approximate embedding can’t be improved upon when embedding a general matroid into a graphic matroid. Can we improve this ratio when embedding into a larger class such as binary matroids:
Open Question 2.
Is there is a sublinear $\beta(n)$ such that any matroid M admits a $\beta(n)$-approximate offline embedding into the complete binary matroid?